OpenStudy (anonymous):

22% of 50 is _____? 100 is 250% of ____?

OpenStudy (anonymous):

22% = 22/100 multiply (22/100) * 50 = 11 100 = (250/100)x solve for x

OpenStudy (anonymous):

I'm confused where you got the 50 from to x it by though

OpenStudy (anonymous):

@Euler271

OpenStudy (anonymous):

22% of 50 = 22% * 50